OpenStudy (anonymous):

how do i simply this expression 4^3(ab)^6 (over) 4 (ab)^2

OpenStudy (anonymous):

\[\frac{4^3(ab)^6}{4(ab)^2}\]???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[=4^2(ab)^4\] \[=16a^4b^4\]

OpenStudy (anonymous):

why do you get 4^2

OpenStudy (anonymous):

\[4^3=4*4*4\] \[\frac{4*4*4}{4}\] \[=4*4\]

OpenStudy (anonymous):

ohhh ok thanks !